A monopolist with the cost function c=½q² faces demand curve q=12-p


A. What will be his equilibrium price and quantity ?


B. If the some reason the firm behaves as if it were in a perfectly competitive idustry, what will equilibrium price and quantity ?


c. How much money will the firm require to forgo monopoly profits and behave competitively instead ?

1
Expert's answer
2022-02-25T10:41:06-0500

A. His equilibrium price and quantity are:

MR = MC,

12 - 2q = q,

q = 4 units,

p = 12 - 4 = 8.

B. If the some reason the firm behaves as if it were in a perfectly competitive idustry, then the equilibrium price will be lower and quantity will be higher.

c. The firm will require to forgo some part of monopoly profits and behave competitively instead.